OpenStudy (jordanloveangel):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (jordanloveangel):

@magan @MarziaJackson$ @BunnyBree @Tazmaniadevil @taylor12344 @Walkerman15 @waterineyes @Whitemonsterbunny17 @eliciagarza605273 @emo__gamer101 @redeemed1312 @Rockiee @THIS,KILLCHIC @Tina_Russo @young.money @itsbribro @iRespect @priecetheveil @alesia/bailey @Angelina_Vella @AJLee5986 @secret66 @sailawaywithme @sandrayanez @Savane0621 @savannah152426 @stellawalkerson @SlowerThePriest @dan815 @danish071996 @daylon @DullJackel09 @Gabylovesyou @Here_to_Help15 @Jamielynne @JaScOk @jordanloveangel @Jhannybean @KaMaddie @KyanTheDoodle @KailaMorgan7777 @lakeykey @love_to_love_you @limelife @ZoMbAsH @Countrygirl23 @Clueless156 @Crissy15 @cheerleading4life<3 @confusegirllll @CoolKatt16 @chrissy_xoxo @Clalgee @ValarieAnne @BabyGirlAj @baking4lyf @Broken_heartedgirl123 @BunnyBree @Natalia.lovs.you @morganchavez @marylou004 @Mesopretty @myininaya @Mimi_x3 @mkerr017 @mcnabkt @...Dimps...

OpenStudy (anonymous):

looks like something i am learning but idk

OpenStudy (jordanloveangel):

and u get someone to help me please

OpenStudy (anonymous):

ask @TheSmartOne

OpenStudy (anonymous):

he knows everything

OpenStudy (jordanloveangel):

lol ikr

OpenStudy (anonymous):

yes lol

OpenStudy (jordanloveangel):

can u try to help me please

OpenStudy (jordanloveangel):

@dumbcow

OpenStudy (anonymous):

im bad at math. and idk why @thesmartone isnt coming hmm hold on let me tell him

OpenStudy (anonymous):

oh hes viewing it wow

OpenStudy (jordanloveangel):

lol @TheSmartOne

OpenStudy (jordanloveangel):

can u help me please

OpenStudy (anonymous):

i messaged him lol

OpenStudy (jordanloveangel):

oh kk

OpenStudy (jordanloveangel):

anything?

OpenStudy (jordanloveangel):

@SolomonZelman

OpenStudy (anonymous):

try asking @Nnesha

OpenStudy (jordanloveangel):

@Nnesha can u help me

OpenStudy (anonymous):

sorry im trying to get help as much as i can :p

TheSmartOne (thesmartone):

I was afk

TheSmartOne (thesmartone):

@jordanloveangel Also do not tag so many people as that is mass tagging and I don't want to see you get suspended.

OpenStudy (anonymous):

yeah thats what you want us to think lol

OpenStudy (jordanloveangel):

sorry but can u help or nahh

TheSmartOne (thesmartone):

What type of math is Just wondering...

OpenStudy (jordanloveangel):

algebra 1

OpenStudy (jordanloveangel):

@TheSmartOne The function would be (x-h)^2=4(a)(x-k)^2 is this the right answer

TheSmartOne (thesmartone):

The standard form is (x - h)2 = 4a (y - k), where the focus is (h, k + a) and the directrix is y = k - a. If the parabola is rotated so that its vertex is (h,k) and its axis of symmetry is parallel to the x-axis, it has an equation of (y - k)2 = 4a (x - h), where the focus is (h + a, k) and the directrix is x = h - a. Got it from : http://jwilson.coe.uga.edu/emt725/class/sarfaty/emt669/instructionalunit/parabolas/parabolas.html

OpenStudy (jordanloveangel):

thx

TheSmartOne (thesmartone):

Is that all you need?

TheSmartOne (thesmartone):

No problem. Sorry for the late reply...